# Keeping in Place After the Storm—Emergency Assistance and Evictions

### v2

@@ Abstract
We offer evidence that federal emergency assistance (FEMA) in the days following natural disasters mitigate evictions in comparison to similar emergency scenarios where FEMA aid is not provided. We find an approximate 16.5% increase in overall evictions after hurricane natural disaster events that increases to 19.7% when excluding ZIP codes that receive FEMA rental assistance. Furthermore, we also show that FEMA aid acts as a liquidity buffer to other forms of emergency credit, specifically we find that both transactions volumes and defaults decrease during hurricane events in locations that do receive FEMA aid. This effect largely reverses in areas that do not receive FEMA aid, where the magnitude of transaction volumes drop by less and default rates remain similar relative to the baseline. Overall, this suggests that the availability of emergency liquidity during natural disaster events is indeed a binding constraint with real household financial consequences, in particular through our documented channel of evictions and in usage of high-cost credit.
@@ 1 Introduction
The loss of structures and in particularly housing is often among the notable and lasting images of in the wake of natural disasters. The resulting financial loss and displacement highlight critical questions over both the individual and government role in pre-disaster mitigation and post-disaster response, in particular through their shortcomings. At the micro level, the loss born by households can be large and has been documented in various forms in the literature.
Historically, the destruction of housing in such events has often led to rebuilt neighborhoods that differ significantly from their original form and community composition. The Boston fire of 1871 (Hornbeck and Keniston, 2017) and the San Francisco 1906 earthquake (Siodla, 2015) provide specific historical moments of migration and reconstruction where entire neighborhoods were rebuilt but also with a notable change in the former population to which (Boustan et al., 2020) find systematic evidence of out-migration following disaster events in the United States. To contrast, in our work, we study instances of assistance in preserving housing access. That is, we examine the existence of federal assistance households to maintain their ability to make housing payments due to potential disruptions and damages directly related to the natural disaster.
Prior to the creation of the Federal Emergency Management Agency under the Carter administration in 1979, many natural disasters that have marked the history of the United States have required federal emergency funding through ad-hoc legislation in the aftermath of their occurrence. One form of disaster relief has been the need for emergency housing in the wake of damage and destruction to housing property. This has further expanded to include rental assistance starting initially under the Stafford Act of 1988 and then further expanded under the Individuals and Households Program in 2002333(Webster, 2022) that aim to recognize disruption in household financial ability due to uninsured or under-insured emergency expenses.
@@ 1.1 Literature Review
The household financial impacts of natural disasters have been increasingly studied, with significant attention given to household resilience and recovery mechanisms. The intersection of disasters and eviction risks has also garnered attention. (Brennan et al., 2022) explore the heightened vulnerability of renters in disaster-affected areas, showing that disasters exacerbate eviction rates, particularly for vulnerable populations. (Raymond et al., 2022) examine the legal and policy dimensions of eviction laws, emphasizing the gaps in tenant protections in states such as Florida. These findings are particularly relevant to our work, underscoring the cascading effects of disasters on eviction, financial distress, and reliance on high-cost credit.
@@
Furthermore, the consequences of hurricane events has notably received attention in the literature. (Gallagher and Hartley, 2017) analyze the financial consequences of hurricane Katrina, highlighting the crucial role of federal aid in stabilizing household finances. (Deryugina et al., 2018) corroborate these findings, using tax return data to show long-term income declines and wealth disruptions. (Del Valle et al., 2024) extend this analysis to hurricane Harvey, examining household decision-making in the aftermath of a disaster. They identify shifts in financial behavior, including increased reliance on savings and external credit. Similarly, (Collier et al., 2024) demonstrates that FEMA assistance mitigates financial distress, using credit files data they show a reduction in delinquencies and defaults in affected regions. Our study complements this result by focusing on evictions and on substitution effects with higher cost credit, particularly how households resort to alternative financial lenders like payday loans when formal aid is absent or insufficient.
@@
The literature on payday loans has highlighted the dual nature of high-cost credit, as both a financial lifeline and a debt trap. We similarly contrast its availability in the context of natural disasters.(Bhutta et al., 2015) describe how payday loans exacerbate financial strain for vulnerable borrowers, while (Morse, 2011) offers a contrasting view, emphasizing their utility during emergencies. (Gathergood et al., 2019) shows that payday loans increase financial distress, especially for low-income borrowers, but regulatory interventions can mitigate harmful effects. In the same vein, (Dobridge, 2016) finds mixed effects, while access to high-cost credit aids short-term liquidity, it negatively affects long-term financial outcomes. Extending this debate, our findings suggest that payday loans act as a substitute for formal disaster relief, especially in regions excluded from FEMA’s coverage. This substitution underscores the dual-edged nature of payday lending in disaster recovery contexts, offering immediate relief but also amplifying financial fragility as we observe a difference in subsequent default rates.
@@
Beyond individual outcomes, the broader economic implications of disasters have been extensively analyzed. (Bernstein et al., 2019) highlight the long-term economic vulnerabilities of disaster-prone areas, particularly the impact of rising sea levels on property prices. The interplay between financial aid and alternative credit sources offers another dimension to understanding recovery dynamics. (Collier et al., 2024) analyze demand for disaster recovery loans, illustrating how credit availability shapes household recovery trajectories. (Lane, 2024) underscores the importance of guaranteed credit programs in building resilience, particularly for lower-income households. (Malmin, 2023) explores the role of federal credit in shaping future wealth trajectories, emphasizing the need for equitable and targeted recovery programs.
@@
Finally, recent studies on disaster aid distribution and its long-term effects highlight systemic inequalities. (Billings et al., 2022) document the inequities in financial aid distribution during Hurricane Harvey, revealing gaps in support for vulnerable populations. (Bufe et al., 2021) explores how financial shocks impact lower-income households, identifying key factors that contribute to financial resilience, such as savings, credit access, and social support networks. (Ratnadiwakara and Venugopal, 2020) finds that flood-prone areas increasingly attract lower-income and less creditworthy populations due to declining property values and housing costs.
@@ 2.1.1 Eviction Data
Princeton’s Eviction Lab data provides Census-Tract level information at a weekly level, including eviction, filing, and at-risk outcome variables Desmond et al. (2018). Covering the period from 2003 to 2017. Table 1 shows summary statistics. The evict variable describes the number of evictions aggregated at the weekly-census tract level. The filing variable describes the number of court filings for eviction and the atrisk variable adjusts this for the number of individuals living in the household at risk of being evicted.
@@ 2.1.3 Hurricane Forecast Cone
The National Hurricane Center (NHC) provides forecast cones for all hurricane and tropical storm events and includes the following characteristics: the forecast date, hurricane category, and the forecast time of impact. We use these cones to construct a hurricane time-of-impact and path dataset by examining all the cones for all hurricane and tropical storm events that traverse the state of Florida
However, the forecast cones are only available as far back as 2008. For storm events prior to 2008, we utilize NHC’s data on the storm’s realized path to reconstruct the forecast cone for each storm using the same properties as the forecast cones provided by the NHC in the period after 2008.444https://www.nhc.noaa.gov/aboutcone.shtml provides details on how the forecast cone is constructed. Figure 1 shows an example of a forecast cone. The first point represents the actual location of the hurricane, and as we move away from the initial point, the cone covers a larger area with increasing uncertainty, a method of construction chosen by the NHC based on the prediction error over the past five years.
@@ 2.2 Methodology
2.2.1 Construction of treatment and control groups
To measure the impact of hurricanes on households, we construct two groups: treatment and control. The treated group consists of ZIP codes (or census tracts) that were within the forecast cone at any point for the entire path of the storm event, from the first prediction to the storm’s dissipation. These areas appear in the forecast cone and are likely to experience some direct or/and indirect effects from the hurricane.
In contrast, our control group includes all ZIP codes (census tracts) that remain completely outside the forecast cone throughout the storm’s progression. This ensures that the behavior and outcomes in these areas were more likely to not be influenced by the forecast or the storm itself, allowing us to isolate the impact of the disaster.
Figures 3 and 3 provide insight into our strategy. They show the evolution of four forecast cones for hurricane Hermine in 2016. The treated group is the green area, whereas the control group is the intersection between all blue areas, as demonstrated in Figure 4.
Figure 2: Forecast cones: distinguish between areas inside and outside the cone
Figure 3: Forecast cones: Evolution of forecast cones during the path of the storm with an example of the treated group (green area)
Figure 4 provides insight into how we constructed the treatment and control groups.
Control group (Blue): This group includes ZIP codes (census tracts) that are entirely outside of any given forecast cone, meaning that these areas were not expected to be affected by a hurricane at any point in time. By focusing on areas not under the forecast at any point, we construct a control group whose behavior and outcomes are assumed to be unaffected directly by storm warnings or the storm itself.
Treated Group (Green): This group includes ZIP codes (census tracts) located within the hurricane forecast cone. However, only areas that remained in the forecast cone up until the closest forecast cone (3 hours prior to impact) are included. That way, we include areas that were eventually either directly in or very close to the realized hurricane path.
In this setup, we compare two groups that likely have very differential impact. We compare the regions that were actually affected (the green area) to the regions (the blue area) that were consistently outside any potential threat of impact. The areas in red were excluded as they were previously in the forecast path but did not eventually remain in the realized path of the storm.
Figure 4: Dynamic 5-Day Forecast Cone
@@ 2.2.2 Event Study Framework
To construct the data for the event study, we treat each hurricane and tropical storm as a separate event study. We merge the constructed treated and control group dataset of storm affected geographies with the evictions dataset at the census tract level for six months before and six months after the event. For the treated group, we use the exact date of the storm path for each census tract and merge with the observed week in the eviction dataset. In the control group, since the geographies were constructed to not intersect with the storm forecast, we assign a pseudo-date, in particular, we use the median storm date. Since our evictions dataset is at the weekly level, we calculate the relative event date in weeks and also in months to the storm date observed. We follow the same approach with payday loans, calculating the relative event date at the daily level since our payday loan data is recorded daily.
@@
As in (Cengiz et al., 2019) and Dube and Lindner (2024), this event study regression aligns events based on event time rather than calendar time, making it comparable to a scenario where all events occur simultaneously. This approach aims to correct well documented challenges in the standard difference-in-differences approach (Baker et al., 2022).
@@ 3.1.1 Hurricanes
Using our evictions dataset we find that ZIP codes that experienced hurricanes saw a monthly average increase of evictions of 0.525 per month per 1000 renters [^5] We normalize the raw number of evictions and filings using census data on the number of renters present in a ZIP code to account for heterogeneity in the percentage of renters and in the population per ZIP code. Table 4. This represents a monthly increase of 16% evictions per month compared to the baseline. The event study in Figure 5 shows that this effect persists up to 6 months after the initial hurricane. The event study also shows a gradual increase in the initial two months up to a peak around months 3-5, reflecting the nature of the eviction process that may lag from the initial missed rental payment date, to the eviction filing, to the finally court approved eviction order.
We also evaluate the change in the number of filings, in which we find an increase of 0.928 in the number of filings per month per 1000 renters, in the 6 month that follow the hurricane event. This represents a 28.7% increase relative to the baseline rate of filings per month.
Figure 5: The figure provides the estimated coefficients from an event study examining the effect of hurricane exposure on monthly eviction-related outcomes at the ZIP code level. The analysis covers a window from five months before to six months after the hurricane event, with the month immediately preceding the event (month -1) serving as the reference period. Panel (a) displays estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates. Vertical lines indicate 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
Table 4: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level. The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes. the coefficient of 0.331 for evictions represents an approximate 16.5% increase relative to the treated group’s pre-event average. Similarly, the 0.428 increase in filings corresponds to about a 13.3% rise from the treated group’s pre-event average. Standard errors clustered at the ZIP code level are reported in square brackets. ***, **, and * indicate statistical significance at the 1%, 5%, and 10% levels, respectively.
Dependent Variable Eviction Filing
Period Monthly Monthly
Window [-5 ; 6] [-5 ; 6]
Treated×$\times$×Post 0.331∗∗∗ 0.428∗∗∗
[0.089] [0.110]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 18501 18501
Table 5: Baseline average monthly level per 1000 renters per ZIP code in the pre-event period.
Variable Control Group Treated Group
Eviction Rate 2.403 2.002
Filing Rate 3.601 3.203
We then investigate the role of FEMA, in particular hurricane events in which FEMA made renter assistance available in impacted ZIP codes. Our results indicate a possible bifurcation in the resultant filings and evictions in the period following the hurricane event. Both the event-study and difference-in-differences regression results illuminate the decrease rate of filings and evictions for hurricane events where FEMA assistance is made available. In particular, we firstly find a decrease in the monthly average of 0.418 and 0.445, for evictions and filings per month, respectively Table 6 when restricting to the event study just within treated units. These represent a 20.9% and 13.9% decrease, respectively, for hurricanes that did not receive FEMA aid. Although both results are not statistically significant, we also show a larger estimate for evictions and filings when excluding ZIP codes that included FEMA assistance in the difference-in-differences results relative to the baseline. Comparatively, evictions increased by a statistically significant 0.392 when excluding FEMA receiving ZIP codes relative to the control group, a larger coefficient than 0.331 result when including all ZIP codes. Similarly, filings increased by a statistically significant 0.502 when excluding FEMA assistance, larger than the 0.428 coefficient when including all ZIP codes.
@@
Figure 6: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly eviction and filing rates at the ZIP code level on the treated. The analysis spans a window from five months before to six months after the hurricane event, with the month preceding the event (month -1) serving as the reference period. Panel (a) presents the interacted monthly eviction estimates for ZIP codes that received FEMA aid versus those that did not for the treated group, while Panel (b) shows monthly interacted filing estimates for ZIP codes that received FEMA aid. Standard errors are clustered at the ZIP code level.
@@
Figure 7: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly eviction and filing rates at the ZIP code level with the entire set of control ZIP codes but excluding the treated ZIP codes that do not receive FEMA assistance. The analysis spans a window from five months before to six months after the hurricane event, with the month preceding the event (month -1) serving as the reference period. Panel (a) presents the interacted monthly eviction estimates for ZIP codes that do not receive FEMA assistance versus the control group, while Panel (b) shows monthly interacted filing estimates for ZIP codes that do not receive FEMA assistance versus the control group. Standard errors are clustered at the ZIP code level.
Dependent Variable Eviction Filing
Period Monthly Monthly
Window [-5 ; 6] [-5 ; 6]
FEMA×$\times$×Post -0.418 -0.445
[0.291] [0.303]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 2729 2729
Table 6: Results for monthly eviction and filing outcomes at the ZIP code level on the treated group only, separated into ZIP codes that received FEMA assistance and those did not. this provides within storm event estimate comparisons for all ZIP codes that were within the storm path. The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes. For FEMA receiving ZIP codes, the coefficient of -0.418, for evictions, while not statistically significant, represents an approximate 20.9% decrease relative to the treated group’s pre-event average (2.002), while the -0.445 decrease in filings corresponds to about a 13.9% decrease from the treated group’s baseline (3.203).
Dependent Variable Eviction Filing
Period Monthly Monthly
Window [-5 ; 6] [-5 ; 6]
Treated×$\times$×Post 0.392∗∗∗ 0.502∗∗∗
[0.092] [0.116]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 16647 16647
Table 7: Difference-in-difference results for monthly eviction and filing outcomes at the ZIP code level while excluding treated ZIP codes that received FEMA assistance. This provides a comparison with control group ZIP codes that for the same time period were not in the storm path. The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes. For non-FEMA receiving ZIP codes, the statistically significant coefficient of 0.392, for evictions, is larger than the 0.331 coefficient for all treated ZIP codes, represents an approximate 19.6% increase relative to the treated group’s pre-event average (2.002), while the 0.502 increase in filings, larger than the 0.428 coefficient for all treated ZIP codes, corresponds to about a 15.6% increase from the treated group’s baseline (3.203).

### v3

@@ Abstract
We offer evidence that federal emergency assistance (FEMA) in the days following natural disasters mitigate evictions in comparison to similar emergency scenarios where FEMA aid is not provided. We find an approximate 10.9% increase in overall evictions after hurricane natural disaster events driven in large part by areas in close proximity of the hurricane path that do not receive FEMA rental assistance. Furthermore, we also show that FEMA aid acts as a liquidity buffer to other forms of emergency credit, specifically we find that both transactions volumes and defaults decrease during hurricane events in locations that do receive FEMA aid. This effect largely reverses in areas that do not receive FEMA aid, where the magnitude of transaction volumes drop by less and default rates remain similar relative to the baseline. Overall, this suggests that the availability of emergency liquidity during natural disaster events is indeed a binding constraint with real household financial consequences, in particular through our documented channel of evictions and in usage of high-cost credit.
@@ 1 Introduction
The loss of structures and in particularly housing is often among the notable and lasting images of in the wake of natural disasters. The resulting financial loss and displacement highlight the importance of both the individual and government role in pre-disaster mitigation and post-disaster response. At the micro level, the loss born by households can be large and has been documented in various forms in the literature.
Historically, the destruction of housing in such events has led to rebuilt neighborhoods that differ significantly from their original form and community composition. The Boston fire of 1871 (hornbeck2017) and the San Francisco 1906 earthquake (SIODLA201548) provide specific historical moments of migration and reconstruction where entire neighborhoods were rebuilt but also with a notable change in the former population to which (BOUSTAN2020103257) find systematic evidence of out-migration following disaster events in the United States. To contrast, in our work, we study instances of assistance in preserving housing access. That is, we examine the existence of federal assistance households to maintain their ability to make housing payments due to potential disruptions and damages directly related to the natural disaster.
Prior to the creation of the Federal Emergency Management Agency under the Carter administration in 1979, many natural disasters that have marked the history of the United States have required federal emergency funding through ad-hoc legislation in the aftermath of their occurrence. One form of disaster relief has been the need for emergency housing in the wake of damage and destruction to housing property. This has further expanded to include rental assistance starting initially under the Stafford Act of 1988 and then further expanded under the Individuals and Households Program in 2002444(crs_r47015) that aim to recognize disruption in household financial ability due to uninsured or under-insured emergency expenses.
@@ 1.1 Literature Review
The household financial impacts of natural disasters have been increasingly studied, with significant attention given to household resilience and recovery mechanisms. The intersection of disasters and eviction risks has also garnered attention. (brennan2022perfect) explore the heightened vulnerability of renters in disaster-affected areas, showing that disasters exacerbate eviction rates, particularly for vulnerable populations. (raymond2022preventing) examine the legal and policy dimensions of eviction laws, emphasizing the gaps in tenant protections in states such as Florida. These findings are particularly relevant to our work, underscoring the cascading effects of disasters on eviction, financial distress, and reliance on high-cost credit.
@@
Furthermore, the consequences of hurricane events has notably received attention in the literature. (gallagher2017household) analyze the financial consequences of hurricane Katrina, highlighting the crucial role of federal aid in stabilizing household finances. (deryugina2018economic) corroborate these findings, using tax return data to show long-term income declines and wealth disruptions. (del2024household) extend this analysis to hurricane Harvey, examining household decision-making in the aftermath of a disaster. They identify shifts in financial behavior, including increased reliance on savings and external credit. Similarly, (collier2024credit) demonstrates that FEMA assistance mitigates financial distress, using credit files data they show a reduction in delinquencies and defaults in affected regions. Our study complements this result by focusing on evictions and on substitution effects with higher cost credit, particularly how households resort to alternative financial lenders like payday loans when formal aid is absent or insufficient.
@@
The literature on payday loans has highlighted the dual nature of high-cost credit, as both a financial lifeline and a debt trap. We similarly contrast its availability in the context of natural disasters.(bhutta2015payday) describe how payday loans exacerbate financial strain for vulnerable borrowers, while (morse2011payday) offers a contrasting view, emphasizing their utility during emergencies. (gathergood2019payday) shows that payday loans increase financial distress, especially for low-income borrowers, but regulatory interventions can mitigate harmful effects. In the same vein, (dobridge2016better) finds mixed effects, while access to high-cost credit aids short-term liquidity, it negatively affects long-term financial outcomes. Extending this debate, our findings suggest that payday loans act as a substitute for formal disaster relief, especially in regions excluded from FEMA’s coverage. This substitution underscores the dual-edged nature of payday lending in disaster recovery contexts, offering immediate relief but also amplifying financial fragility as we observe a difference in subsequent default rates.
@@
Beyond individual outcomes, the broader economic implications of disasters have been extensively analyzed. (bernstein2019disaster) highlight the long-term economic vulnerabilities of disaster-prone areas, particularly the impact of rising sea levels on property prices. The interplay between financial aid and alternative credit sources offers another dimension to understanding recovery dynamics. (collier2024credit) analyze demand for disaster recovery loans, illustrating how credit availability shapes household recovery trajectories. (lane2024adapting) underscores the importance of guaranteed credit programs in building resilience, particularly for lower-income households. (malmin2023access) explores the role of federal credit in shaping future wealth trajectories, emphasizing the need for equitable and targeted recovery programs.
@@
Finally, recent studies on disaster aid distribution and its long-term effects highlight systemic inequalities. (billings2022let) document the inequities in financial aid distribution during Hurricane Harvey, revealing gaps in support for vulnerable populations. (bufe2021financial) explores how financial shocks impact lower-income households, identifying key factors that contribute to financial resilience, such as savings, credit access, and social support networks. (ratnadiwakara2020areas) finds that flood-prone areas increasingly attract lower-income and less creditworthy populations due to declining property values and housing costs.
@@ 2.1.1 Eviction Data
Princeton’s Eviction Lab data provides Census-Tract level information at a weekly level, including eviction, filing, and at-risk outcome variables evictionlab. Covering the period from 2003 to 2017. Table 1 shows summary statistics. The evict variable describes the number of evictions aggregated at the weekly-census tract level. The filing variable describes the number of court filings for eviction and the atrisk variable adjusts this for the number of individuals living in the household at risk of being evicted.
@@ 2.1.3 Hurricane Forecast Cone
The National Hurricane Center (NHC) provides authoritative information on the timing, location, and progression of tropical storms and hurricanes affecting the United States. In our study, this data is critical for reconstructing the geographic footprint of storms that impacted Florida between 2004 and 2018. However, the forecast cones are only available as far back as 2008.
For storm events prior to 2008, we relied on NHC forecast cone archives, [^5] https://www.nhc.noaa.gov/aboutcone.shtml provides details on how the forecast cone is constructed. which record the storm’s position and projected path at successive forecast times. From these records, we extracted the initial forecast points that represent the best estimate of the storm’s current location, thereby tracing the realized path of each hurricane. Figure 1 shows an example of a forecast cone. The first point represents the actual location of the hurricane, and as we move away from the initial point, the cone covers a larger area with increasing uncertainty, a method of construction chosen by the NHC based on the prediction error over the past five years.
For storm events prior to 2008, we complemented NHC records with storm track data from NOAA’s Historical Hurricane Tracks database. By combining these sources, we were able to build a consistent dataset of storm paths across the full study period. This provided the foundation for identifying geographic areas affected by each hurricane and for distinguishing treated from control communities in the empirical analysis.
@@ 2.2 Methodology
2.2.1 Construction of treatment and control group
To assess the impact of hurricanes on households, we classify communities into treatment and control groups based on their geographic proximity to the storm path. The treated group consists of ZIP codes whose boundaries intersect the realized hurricane track, extended by a buffer. This buffer reflects the forecast uncertainty at the initial advisory position, as defined by the National Hurricane Center (NHC) when constructing its cones of uncertainty. It captures not only areas directly along the storm’s path but also nearby regions likely to experience significant effects. In this way, the treated group represents communities most plausibly exposed to wind, rainfall, and other storm-related disruptions.
By contrast, the control group includes ZIP codes in Florida that did not intersect the buffered storm path. These areas lie outside the spatial footprint of the hurricane and thus provide a counterfactual for how outcomes would have evolved in the absence of storm exposure. This distinction ensures that observed differences between the treated and control groups can be more directly attributed to hurricane impacts rather than broader regional dynamics.
Figure 2 illustrates our spatial treatment assignment strategy using Hurricane Charley (2004) as an example. The map shows how ZIP Code Tabulation Areas (ZCTAs) intersecting the storm’s path are classified as treated, while those outside the exposure zone serve as controls.
Figure 2: Treatment and Control ZIP Codes for Hurricane Charley (2004). The figure shows the hurricane’s observed track (black points), the buffer zone used to define exposure (light blue), and ZIP Codes classified as treated (red) or control (blue).
@@ 2.2.2 Event Study Framework
To construct the data for the event study, we treat each hurricane and tropical storm as a separate event study. We merge the constructed treated and control group dataset of storm affected geographies with the evictions dataset at the ZIP code level for six months before and six months after the event. For the treated group, we use the exact date of the storm path for each ZIP code and merge with the observed week in the eviction dataset. For the control group, whose geographies were constructed to remain outside the storm’s path, we assign a pseudo-date: when the storm event occurred on a single day, we assign that same day to all control ZIP codes, whereas for multi-day events, we randomly distribute the storm dates across control ZIP codes (e.g., if the storm lasted two days, some control ZIP codes are assigned the first day while others are assigned the second). Since our evictions dataset is at the weekly level, we calculate the relative event date in weeks and also in months to the storm date observed. We follow the same approach with payday loans, calculating the relative event date at the daily level since our payday loan data is recorded daily.
@@
As in (cengiz) and dube2024minimum, this event study regression aligns events based on event time rather than calendar time, making it comparable to a scenario where all events occur simultaneously. This approach aims to correct well documented challenges in the standard difference-in-differences approach (Baker).
@@ 3.1.1 Hurricanes
Using our evictions dataset we find that ZIP codes that experienced hurricanes saw a maximum monthly average increase of evictions of 0.201 per month per 1000 renters [^6] We normalize the raw number of evictions and filings using census data on the number of renters present in a ZIP code to account for heterogeneity in the percentage of renters and in the population per ZIP code. as in Table 4. This represents a monthly increase of 10.9% evictions per month compared to the baseline. The event study in Figure 3 shows that this effect persists up to 6 months after the initial hurricane. The event study also shows a increase around months 2-5, reflecting the nature of the eviction process that may lag from the initial missed rental payment date, to the eviction filing, to the finally court approved eviction order.
We also show in Table 4 that including ZIP codes by broadening the size of the treatment group radius (in nautical miles) results in a decreasing effect size for evictions, as one might expect given the damage from the hurricane event correlates with distance. At 100 nautical miles the effect essentially disappears. This result is important as we later show that this effect does not hold when separating into FEMA receiving and non-FEMA receiving ZIP codes.
@@
Figure 3: Estimated coefficients from an event study model evaluating the impact of hurricane exposure on monthly eviction and filing rates at the ZIP code level with a treatment radius of 21 nautical miles. The event window spans from five months before to six months after landfall, with the month immediately preceding the hurricane (month -1) serving as the reference period. Standard errors are clustered at the ZIP code level.
@@
Figure 4: Estimated coefficients from an event study model evaluating the impact of hurricane exposure on monthly eviction and filing rates at the ZIP code level with a treatment radius of 50 nautical miles. The event window spans from five months before to six months after landfall, with the month immediately preceding the hurricane (month -1) serving as the reference period. Standard errors are clustered at the ZIP code level.
Treatment Radius
10 15 21 26 50 100
Panel A: Eviction
[…]
[0.100] [0.097] [0.103] [0.114] [0.114] [0.084]
Panel B: Filing
[…]
[0.132] [0.126] [0.125] [0.137] [0.150] [0.149]
Event-specific ZIP Code FE Y Y Y Y Y Y
Event-specific Time FE Y Y Y Y Y Y
Observations 22,638 21,617 20,383 19,360 16,939 15,894
Table 4: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level by treatment radius. The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes. Standard errors clustered at the ZIP code level are reported in square brackets. ***, **, and * indicate statistical significance at the 1%, 5%, and 10% levels, respectively.
Treatment Radius
10 15 21 26 50 100
Panel A: Evictions
Control 2.096 2.112 2.189 2.239 2.302 2.201
Treated 1.956 1.974 1.842 1.782 1.782 1.824
Panel B: Filings
Control 3.129 3.164 3.291 3.372 3.515 3.535
Treated 3.058 3.041 2.811 2.740 2.746 2.693
Table 5: Baseline averages by group (pre-hurricane period) for each treatment radius. Control and treated ZIP codes are compared across radii of 10, 15, 21, 26, 50, and 100 miles.
Figure 5: A bin scatter plot of the proportion of ZIP codes at increasing distances from the hurricane path that had recipients of FEMA rental assistance. The plot shows a decreasing in the proportion with a notable jump at the 20 nautical mile mark.
We then investigate the role of FEMA, in particular hurricane events in which FEMA made renter assistance available in impacted ZIP codes. Eligibility for assistance through FEMA depends on a geography being declared FEMA eligible [^7] FEMA provides a web page for individuals to lookup whether their ZIP code lies in a FEMA disaster declared county https://www.fema.gov/node/how-do-i-know-if-my-area-eligible-assistance?, as well as individual assessments of damage and hardship which correlate with proximity to the path of the hurricane event. Figure 6 shows the ZIP codes that received assistance via FEMA and their distance to the hurricane event. We observe that the intersection of ZIP code boundaries, storm damage intensity and receiving FEMA assistance correlate and vary in relation to distance. Figure 5 shows the proportion of ZIP codes with individuals receiving FEMA assistance and the distance from the path of the hurricane during FEMA declared hurricanes.
Figure 6: A map of ZIP codes that received FEMA rental assistance in red during Hurricane Charley. The orange non-FEMA treated ZIP codes are those that fall within the hurricane cone but did not receive FEMA rental assistance. The map displays two cone sizes, 21 nautical miles and 50 nautical miles, illustrating the sensitivity of the proportion of ZIP codes receiving FEMA rental assistance at different distances.
Building on our baseline eviction and filings results we observe a bifurcation in the resultant filings and evictions in the period following the hurricane event that varies with FEMA assistance. Both the event-study and difference-in-differences regression results illuminate the decrease rate of filings and evictions for hurricane events where FEMA assistance is made available. In particular, from Table 6 we find a maximal bifurcation in the monthly average evictions with an increase of 0.561 in non-FEMA receiving ZIP codes and a 0.541 increase in filings at the 50 nautical mile radius of inclusion of ZIP codes.
Figure 7: Estimated coefficients from an event study model evaluating the impact of hurricane exposure on monthly eviction rates at the ZIP code level with a treatment radius of 21 nautical miles. The event window spans from five months before to six months after landfall, with the month immediately preceding the hurricane (month -1) serving as the reference period. The treated group is divided based on FEMA rental assistance availability. Standard errors are clustered at the ZIP code level.
Figure 8: Estimated coefficients from an event study model evaluating the impact of hurricane exposure on monthly eviction rates at the ZIP code level with a treatment radius of 50 nautical miles. The event window spans from five months before to six months after landfall, with the month immediately preceding the hurricane (month -1) serving as the reference period. The treated group is divided based on FEMA rental assistance availability. Standard errors are clustered at the ZIP code level.
Treatment Radius
10 15 21 26 50 100
Panel A: Eviction
[…]
[0.159] [0.156] [0.164] [0.164] [0.218] [0.141]
[…]
[0.106] [0.099] [0.100] [0.111] [0.107] [0.088]
Panel B: Filing
[…]
[0.162] [0.174] [0.186] [0.217] [0.247] [0.215]
[…]
[0.140] [0.132] [0.125] [0.131] [0.148] [0.146]
Event-specific ZIP Code FE Y Y Y Y Y Y
Event-specific Time FE Y Y Y Y Y Y
Observations 22,638 21,617 20,383 19,360 16,939 15,894
Table 6: Difference-in-Differences estimates of post-event eviction and filing outcomes by FEMA-designated and non-FEMA-designated areas across treatment radii. Standard errors clustered at the ZIP code level are in brackets. ***, **, and * indicate significance at the 1%, 5%, and 10% levels, respectively.